OpenStudy (anonymous):

What is the point-slope form of a line with slope 2/5 that contains the point (-3, 6)?

OpenStudy (noseboy908):

point-slope form is "y-y1=m(x-x1)" simply put the points and slope into the equation and solve for y.

OpenStudy (anonymous):

so do you know the answer?

OpenStudy (noseboy908):

yes, but what do you think it is?

OpenStudy (anonymous):

y - 6 = -(x+ 3) i think this is the answers

OpenStudy (anonymous):

is it?

OpenStudy (noseboy908):

except for the slope, that step is correct.

OpenStudy (noseboy908):

m =2/5, not -1.

OpenStudy (anonymous):

ooo ok thank you!!!:)

OpenStudy (noseboy908):

you're welcome. make sure to hit best response!

OpenStudy (noseboy908):

thank you.

OpenStudy (anonymous):

your welcome

OpenStudy (noseboy908):

*you're

OpenStudy (anonymous):

what?

OpenStudy (noseboy908):

nothing. carry on.